OpenStudy (anonymous):

The lengths of the sides of a right triangle are a and b, and the hypotenuse is c. Find the area of the triangle. b = 2 in.; c = 6 in. A = sq. in.

OpenStudy (anonymous):



OpenStudy (lgbasallote):

a would be the base b would be the height

OpenStudy (anonymous):

thats correct

OpenStudy (lgbasallote):

you can solve for a by doing \[\large \sqrt{6^2 - 2^2}\]

OpenStudy (anonymous):

area = 1/2(bh), in your case area = 1/2(ab) b = 2 a = 36 - 4 = 32. \[\sqrt{32} = 4\sqrt{2}. A = 4\sqrt{2}\] area = 1/2(2)(4sqrt{2}) area = 4sqrt{2}

OpenStudy (lgbasallote):

\[a = \sqrt{36 - 4} = \sqrt{32} = 4\sqrt 2\]

OpenStudy (lgbasallote):

use formula of area \[\large A = \frac{4 \sqrt 2 \times 2}{2}\]

OpenStudy (anonymous):

I also have this problem. The lengths of the sides of a right triangle are a and b, and the hypotenuse is c. Find the area of the triangle. b = 2 in.; c = 6 in. A = sq. in.

OpenStudy (lgbasallote):

i just answered the question

OpenStudy (lgbasallote):

what i wrote is easy enough to simplify... \[\Large \frac{4\sqrt 2 \times 2}{2}\]

OpenStudy (anonymous):

sorry it didn't type right ....The lengths of the sides of a right triangle are a and b, and the hypotenuse is c. Find the area of the triangle. A = 2/3 ft.; c = 4 ft. A = sq. ft.